OpenStudy (anonymous):

Re-write the expression as an algebraic expression in x and y (that is, with no trigonometric functions):

OpenStudy (anonymous):

\[\tan (\sin^{-1} x+\cos^{-1} x)\]

OpenStudy (anonymous):

When I calculate that I get an undefined expression. Since your question mentions some variable y, I have to ask if you meant one of those x's to be a y.

OpenStudy (anonymous):

Sorry, you're right. It should be a y instead of an x in cos^-1.

OpenStudy (anonymous):

\[ \tan(\sin^{-1}x + \cos^{-1}y) = \frac{\sin(\sin^{-1}x + \cos^{-1} y)}{\cos(\sin^{-1}x + \cos^{-1}y)} \] \[\sin(\sin^{-1}x + \cos^{-1}y) = \sin(\sin^{-1}x)\cos(\cos^{-1}y) + \sin(\cos^{-1}y)\cos(\sin^{-1}x)\] \[ = xy + \sqrt{1-\cos^2(\cos^{-1}y)}\sqrt{1-\sin^2(\sin^{-1}x)} = xy + \sqrt{1-y^2}\sqrt{1-x^2}\] simiarly, the cosines can be rewritten to yield \[ \cos(\sin^{-1}x + \cos^{-1}y) = \sqrt{1-x^2}\cdot y - x\cdot \sqrt{1-y^2}\]put together, that yields \[ \frac{xy + \sqrt{1-x^2}\sqrt{1-y^2}}{y\sqrt{1-x^2} - x\sqrt{1-y^2}} \]